Solve the inequality log2 (x2 – 3x + 4) <1. Give the answer as an interval.
Answer #1

1. 109 (9.31 3 ห +1) 2.1 -) (- 21 + 1): 2 -)(**.2ฯ - วฯ +4 -2) : 0 (n-3n +2) (*-21 1+2) = 0 * (ฯ - 2) (1)</ (1-2) = 0 (-1) -